OpenStudy (yacoub1993):

The Fiedler family has up to $130,000 to invest. They decide that they want to have at least $40,000 invested in stable bonds yielding 5.5% and that no more than $60,000 should be invested in more volatile bonds yielding 11%. How much should they invest in each type of bond to maximize income if the amount in the stable bond should not exceed the amount in the more volatile bond? What is the maximum income?

OpenStudy (yacoub1993):

@TheRealMeeeee

OpenStudy (therealmeeeee):

first define your variables let x = amount of money (in dollars) invested in the stable bonds at 5.5% y = amount of money (in dollars) invested in the volatile bonds at 11%

OpenStudy (therealmeeeee):

They have up to $130,000 to invest, so this means that the total they have to invest cannot be larger than 130,00 so this means total <= 130000 x+y <= 130000 since x+y is the total amount to invest (in both combined)

OpenStudy (yacoub1993):

then

OpenStudy (therealmeeeee):

we're also told that "they want to have at least $40,000 invested in stable bonds yielding 5.5%", so we know x >= 40000 and we're given that "no more than $60,000 should be invested in more volatile bonds yielding 11%", so we can say y <= 60000

OpenStudy (therealmeeeee):

Income is maximized when x,y are largest possible given restraints, this is at point (60,000, 60,000)

OpenStudy (therealmeeeee):

In the volatile bonds = not > than 60,000 $ To get maximum yield , you have to invest 60,000 $ in 11% bonds So amount invested in stable bonds = 130.000 - 60,000 = 70,000 $ This is more than 60,000 $ Both the conditions given cannot be satisfied

OpenStudy (therealmeeeee):

so there ya go